Given a population of the scores of BSHS honor students in Mathematics as 1, 3, 4, 6, and 8. Suppose a sample size of 3 have to be drawn from it for the first, second and third rank;


1
Expert's answer
2021-11-09T16:54:30-0500

a.

We have population values "1,3,4,6,8," population size "N=5"

"\mu=\dfrac{1+3+4+6+8}{5}=4.4"

b.



"\sigma^2=\dfrac{1}{5}((1-4.4)^2+(3-4.4)^2+(4-4.4)^2""+(6-4.4)^2)+(8-4.4)^2)=5.84"


c.



"\sigma=\sqrt{\sigma^2}=\sqrt{5.84}\approx2.4166"


d. We have population values "1,3,4,6,8" population size "N=5" and sample size "n=3." Thus, the number of possible samples which can be drawn without replacement is




"\dbinom{N}{n}=\dbinom{5}{3}=10"


The sampling distribution of the sample means.


"E(\bar{X})=\sum\bar{X}f(\bar{X})=4.4"


The mean of the sampling distribution of the sample means is equal to the the mean of the population.



"E(\bar{X})=\mu_{\bar{X}}=4.4=\mu"



e.



"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"





"=\dfrac{61}{3}-(4.4)^2=\dfrac{2.92}{3}"





"\sigma_{\bar{X}}=\sqrt{Var(\bar{X})}=\sqrt{\dfrac{2.92}{3}}\approx0.9866"

Verification:



"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{5.84}{3}(\dfrac{5-3}{5-1})"





"=\dfrac{2.92}{3}, True"